# Learning to combine top-down context and feed-forward representations under ambiguity with apical and basal dendrites

### v1

@@ 2.3 Single neuron mechanisms of top-down modulation
We next sought to understand the computational role played by top-down modulation of pyramidal apical dendrites in solving this contextual integration task. To this end, we developed a method to identify neurons relevant for solving the task. We then analysed their activations across the various contextualized input conditions to reveal the differential role played by their apical compartments. First, to define the sub-populations of pyramidal neurons performing specific computations, we employed layer-wise relevance propagation (LRP; [57]; supplementary Fig. 1a), a method typically used to explain which regions of an input contribute most to model output. Specifically, each neuron receives a score for its contribution to the model output for each image (Supplementary Fig. 1b). We then defined the subset of neurons most relevant for identifying a given class, $S_{C}$, as the minimum set of neurons required to account for 95% of total neuronal relevance when processing images from that class (Fig.  1c,d, see Methods). Sets of relevant neurons defined in this way exhibit low overlap between classes (Supplementary Fig. 1b-d, see Methods). We observed similar representational separability between classes for other scenario and given appropriate top-down inputs (Supplementary Fig. 1e). Additionally, these sets of relevant neurons are typically sparse and represent $\sim$15of the total neuronal population(Supplementary Fig. 1b,c).
[…]
@@
To evaluate this hypothesis, we applied a mask on selected neurons at inference time and assessed the model’s performance. Specifically, we assess readout test set accuracy when inputs are ambiguous and context is helpful, and compare cases where class-pair overlapping neurons (as identified with LRP) are specifically masked (by setting their value to 0) to cases where other randomly selected neurons are masked. We also evaluate the effect of mask size and found that masking class-pair overlapping neurons specifically lead to significantly more rapid degradation of accuracy with increasing mask size compared to other neurons (Fig. 4d). This highlights the key role of these neurons’ apical compartments in integrating relevant contextual signals.
@@
Figure 4: Class-pair overlapping neurons receive the strongest top-down signals. a, using layer-wise relevance propagation, we identify neurons most relevant for either of two classes (red and blue), or overlap of both classes (magenta). b, amplitude of the apical compartment (normalized over all neurons) for each input/context condition. Three distinct classes of neurons are compared: neurons most relevant for unambiguous class A (blue), B (red), or neurons most relevant for both A and B classes (magenta). c, we hypothesize that class-pair overlapping neurons receive the highest amplitude of top-down signals when processing ambiguous images, especially given matching context.
@@ 2.4 Deriving context from temporal information
Task context ambig.Match unambig.Match unambig.Irrel ambig.Irrel unambig.Contra MNIST oracle 0.996 $\pm$ 0.000 0.983 $\pm$ 0.001 0.979 $\pm$ 0.002 0.461 $\pm$ 0.003 0.964 $\pm$ 0.001 MNIST temporal 0.963 $\pm$ 0.001 0.999 $\pm$ 0.000 0.991 $\pm$ 0.001 0.487 $\pm$ 0.005 0.990 $\pm$ 0.001 EMNIST oracle 0.993 $\pm$ 0.001 0.969 $\pm$ 0.003 0.936 $\pm$ 0.003 0.469 $\pm$ 0.004 0.903 $\pm$ 0.002
Table 1: Readout test set accuracy on MNIST and EMNIST with different context representations. Results are expressed for each scenario as mean % $\pm$ standard deviation.
@@
Table 2: Ablation experiments on EMNIST for the integration rule. Context only $=0$ when the topdown network has access to feedforward latent $\mathbf{Ub}$ and $c$ (standard case), and context only $=1$ when only $c$ is provided as input. Three modulation (integration) rules are tested: hadamard (ours), sum, and concat (see 3. Concat refers to a concatenation of ’topdown’ weights on the basal weight matrix, akin to adding dendritic branches on the basal compartment.
ambig.Match unambig.Match unambig.Irrel ambig.Irrel unambig.Contra modulation context_only
0.989 $\pm$ 0.002 0.959 $\pm$ 0.003 0.931 $\pm$ 0.004 0.469 $\pm$ 0.003 0.841 $\pm$ 0.004 concat 0
0.794 $\pm$ 0.005 0.967 $\pm$ 0.003 0.938 $\pm$ 0.003 0.473 $\pm$ 0.004 0.903 $\pm$ 0.014 concat 1
0.993 $\pm$ 0.001 0.969 $\pm$ 0.003 0.936 $\pm$ 0.003 0.469 $\pm$ 0.004 0.903 $\pm$ 0.001 hadamard 0
0.913 $\pm$ 0.008 0.929 $\pm$ 0.004 0.901 $\pm$ 0.006 0.468 $\pm$ 0.005 0.738 $\pm$ 0.008 hadamard 1
0.994 $\pm$ 0.001 0.974 $\pm$ 0.001 0.944 $\pm$ 0.001 0.466 $\pm$ 0.007 0.920 $\pm$ 0.001 sum 0
0.792 $\pm$ 0.011 0.967 $\pm$ 0.005 0.937 $\pm$ 0.003 0.462 $\pm$ 0.013 0.904 $\pm$ 0.008 sum 1
@@ 3 Discussion
The exact mechanism by which top-down signals exert their functional role in pyramidal neurons remains unknown. One hypothesis is that, in the neocortex, apical dendrites of pyramidal neurons integrate top-down information from higher-order regions, which in turn modulate somatic firing rate for accurate and robust perception. A computational framework that respects known biophysical constraints while leveraging a task that requires integrating different types of context is ideal for testing this hypothesis. Here, we present (1) a functional model of contextual integration at pyramidal neurons that relies on the biophysical properties of apical and basal dendrites, and (2) a novel image classification task and dataset using a conditional generative model. Specifically for (1), we implement a Hadamard (multiplicative elementwise) integration rule based on known biophysical properties of apical dendrites which act as gain modulators of somatic activity [105, 80]. To solve (2), we train the top-down network (mapping of context onto apical dendrites) with the proposed loss function 5 by gradient descent, and assess the role of top-down signals in solving the task across 5 distinct input/context conditions. We find that the top-down network learns to refine ambiguous representations by applying sparse gain modulation to basal signals when contextual information is both available and relevant, but also learns to ignore contradictory or irrelevant contexts. This sparse selectivity is consistent with biased competition in the top-down modulation of pyramidal neurons. Our results show that it is in principle possible for local pyramidal circuits to learn a mapping of contextual representations onto their apical compartments to resolve perceptual ambiguity. Further analysis reveals that a sparse subset of class-pair overlapping neurons are preferentially gain-modulated by top-down contextual inputs, providing a candidate mechanism for how the model solves the perceptual ambiguities. This mixed selectivity could present distinct computational advantages across changing contexts [95].
3.1 Roles of top-down modulation
Top-down feedback is a prominent feature of the neocortical architecture for which several roles have been hypothesized. During learning, it has been proposed to act as a credit assignment signal [81, 70, 86, 59] and/or representing predictions [94, 93]. Besides its role in credit assignment, top-down feedback has also been proposed to support attentional processes [82] by dynamically highlighting the most relevant parts of an input, in both biological [100] and artificial systems [87], and could augment sensory processing with robustness to adversarial stimuli [61] or integration of contextual or task information [91, 106] . Given their biophysical properties, apical dendrites are particularly well poised to integrate these top-down contextual signals, and from a computational perspective, have been shown to be useful for multitask learning [106]. Our computational framework falls into the category of updating feedforward representations. We do not exclude other hypothesized roles for top-down feedback, however their unification under one framework is outside of the scope of this work.
Extending previous work on integration of topdown contextual signals via apical and basal dendrites, we focus specifically on a task in which the model must learn to differentially integrate relevant and irrelevant contextual signals at apical dendrites for accurate perception. This reflects the reality for brains, where not all contextual signals are useful or relevant to the task at hand.
3.2 Role of dendritic integration algorithms on function
Previous models have proposed to integrate feed-forward and feedback information streams using additive [87, 61] , multiplicative [91] rules, or both [106]. In [87, ], top-down signals come from past temporal context from higher-order modules, and are integrated with the feedforward stream via additive key-value attention ([103]). These works explore similar integration rules to ours, but the mathematical forms are not equivalent. Here we implement a Hadamard integration rule on the basis of a direct correspondence to biophysical principles. While the contextual integration task can be solved using different rules 2, under uncertainty, Hadamard integration enables the model to ignore contradictory contextual signals more efficiently compared to additive integration. Furthermore, as basal and apical signals are combined post-activation, our rule can only amplify already active neurons, unlike [106] and [91]. From an efficiency perspective, this eliminates the need to drive other neurons to fire, and uses less of the network capacity, at the cost of expressivity. It remains to be demonstrated whether our rule enables the model to perform well for more complex tasks.
3.3 Testable predictions
We find that our model can learn a non-linear mapping of top-down signals onto the apical compartment, a property that was recently described in pyramidal neurons of the visual cortex [67]. In addition to this compatibility with in vivo conditions, we propose several key links between experimental work and our model. Firstly, we found that under ambiguity, our model displayed a high diversity of neural responses as well as neurons that exhibit highly overlapping responses for specific class-pairs similar to in vivo reports that subsets of pyramidal neurons display a variety of responses given a mismatch between context and sensory evidence [74, 95]. Secondly, we find that masking apical compartments of class-pair overlapping neurons decreased task performance by a significant margin, up to nearly chance level. This is in contrast with in vivo experiments in non-human primates showing that inhibiting higher-order regions, and effectively shifting visual cortex representations to a strictly feed-forward regime, increased performance delays but did not prevent object recognition even for challenging images [75]. Finally, our model could learn a mapping of top-down feedback to the apical compartment independent of learning contextual representations which is compatible with high-order cortical regions, including the hippocampus, learning and updating contextual representations in absence of sensory inputs and yet play a major role in memory retrieval that necessitate contextual information [84].
One key testable prediction from our model is that neurons displaying mixed-selectivity would receive the strongest top-down signals at their apical dendrites. Testing this would require first conducting a controlled contextual integration task with an animal, and identifying a population from which a high degree of overlap is observed in neural responses for multiple categories or task variables. We hypothesize that silencing the apical compartments of these neurons would impair the ambiguity resolution ability of the animal.
3.4 Limitations and outlook
@@
While our model combines functionality with key biophysical properties of dendrites, there are some important assumptions and simplifications that need to be considered. Firstly, our model is designed to implement some key features of cortical pyramidal neurons without focusing on a specific subpopulation that make up cortical layers, including layer 2/3 and layer 5 pyramidal neurons. Here, we assume that contextual signals are provided by a neurons in a higher-order region, which would likely correspond to layer 5 output, and integrated in apical dendrites of putative layer 2/3 pyramidal neurons [58, 96]. It is noteworthy though that layer 5 neurons can also send top-down projections to the basal compartment of layer 2/3 neurons, although to a lesser extent [72].
Secondly, we use a deep learning framework that does not respect Dale’s principle: namely, synaptic weights can be either negative and positive, giving the ability of presynaptic neurons to be both excitatory and inhibitory at the same time for distinct postsynaptic neurons. Although our model ignores this principle, we apply normalization across inputs, which could in principle mimic a role of the inhibitory neurons if they are non-selective. Additionally, Dale’s principle could theoretically be implemented to our deep learning framework with little sacrifice in performance [62].
@@
Our model heavily abstracts temporal dynamics that are normally observed in biological circuits. Specifically, perception is associated with ever evolving neurophysiological responses that anchor to discrete states of sensory experience. For example, in the presence of ambiguous stimuli, human subjects can oscillate between two interpretations, a phenomenon termed ’perceptual bistability’ [71]. During perceptual bistability, neural activity alternates between discrete states [54, 55], following principles of continuous attractor dynamics [63]. In the future, our model could support more realistic temporal dynamics, which would enable the testing of a rich array of hypotheses regarding contextual integration.
In the future, we propose that our model could be extended to hierarchies of sensory modalities interacting with each other. Deep learning models that implement multi-modal interactions are an active area of research [92, 56, 89]. In these cases, representations from each modality are mapped to a joint high dimensional space and optimized via a contrastive self-supervised training objective. These models do not have the explicit objective to resolve conflicting or contradictory information from other modalities. In this context, our framework could be extended to multi-modal self-supervised learning for more robust perception in deep neural networks. Finally, such a model could be leveraged to test specific hypotheses of contextual integration in health and disease, and in particular in cases where the physiological properties of pyramidal neurons are known to be altered [76].

### v2

@@ 2.3 Single neuron mechanisms of top-down modulation
We next sought to understand the computational role of top-down modulation of pyramidal apical dendrites in solving this contextual integration task. To this end, we used a method from the explainable AI literature to rank and identify the neurons most relevant for solving the task. We then compared their activations across the various context-input scenarios. First, to define the most relevant neurons, we employed layer-wise relevance propagation (LRP; (Bach et al., 2015); supplementary Fig. 1a), a common method in explainable AI literature to explain, at each layer, the input’s contribution to the model output. Each neuron receives a score for its contribution to the model output for each image (Supplementary Fig. 1b). We then defined the subset of neurons most relevant for identifying a given class, $S_{C}$, as the minimum set of neurons required to account for 95% of total neuronal relevance when processing images from that class (Fig.  1c,d, see Methods). Sets of relevant neurons defined in this way exhibit low overlap between classes (Supplementary Fig. 1b-d, see Methods). We observed similar representational separability between classes for other scenarios and given appropriate top-down inputs (Supplementary Fig. 1e). Additionally, these sets of relevant neurons are typically sparse and represent $\sim$15of the total neuronal population (Supplementary Fig. 1b,c).
It is noteworthy that our model is not explicitly provided with the ambiguous nature of input images, and therefore learns to extract this information to solve the task. In addition, the model learns to ignore top-down signals when the image is unambiguous and context is contradictory or irrelevant. Given that the apical compartment must combine contextual information and the contextually naive $\mu$ representation differentially based on input ambiguity, we hypothesized that solving the task effectively required the gain modulation of a subset of neurons encoding features specific to the relevant class. To assess this possibility, we registered the amplitude of apical signals arriving at LRP subsets pertaining to class-specific neurons and class-pair overlapping neurons in each scenario (Fig. 4a-f). When comparing all input/context scenarios, we found that apical signals were highest when input images were ambiguous and context informative (1-ANOVA, F416=237.024, p $<$ 0.0001, for the main effect of input/context scenario; Fig. 4b). In the ambig.Match scenario, neurons most relevant for the target class were associated with the highest amplitude in the apical compartment (1-ANOVA, F417=17.9445, p $<$ 0.0001, for the main effect of neuron class; Fig. 4b). We hypothesized that target class specific neurons are also most modulated by top-down context, and that modulating these neurons is necessary to solve the perceptual ambiguity classification task.(Fig. 4c).
@@
To evaluate this hypothesis, we applied a mask on selected neurons post-training, and assessed the model’s ability to solve the task. Specifically, we compare test set accuracy when inputs are ambiguous and context is relevant, under the condition where neurons (as identified with LRP) encoding features specific to the relevant class (which we refer to as context-relevant neurons) are masked out, compared to the condition where random neurons are masked. We also evaluate the effect of mask size and found that masking the context-relevant neurons specifically led to a steeper degradation of accuracy as we increase the number of masked neurons. (Fig. 4d). This highlights the key role of these neurons’ apical compartments in integrating relevant contextual signals.
@@
Figure 4: Top-down has information about ambiguity, and in cases where inputs are ambiguous and context is aligned, resolves ambiguity by amplifying context-relevant neurons.. a, using layer-wise relevance propagation, we identify neurons most relevant for either of two classes (red and blue), or overlap of both classes (magenta). b, amplitude of the apical compartment (normalized over all neurons) for each input/context condition. Three distinct classes of neurons are compared: neurons most relevant for unambiguous class A (blue), B (red), or neurons most relevant for both A and B classes (magenta). c, we hypothesize that context-relevant neurons receive the highest amplitude of top-down signals when processing ambiguous images, especially given matching context.
@@ 2.4 Deriving context from temporal information
Task Context signal ambig.Match unambig.Match unambig.Irrel ambig.Irrel unambig.Contra seq-aMNIST temporal 96.3 $\pm$ 0.1 99.9 $\pm$ 0.0 99.1 $\pm$ 0.1 48.7 $\pm$ 0.5 99.0 $\pm$ 0.1
Table 2: Sequential aMNIST (seq-aMNIST) test set classification accuracy across input-context scenarios with the hadamard integration rule. Results are expressed for each scenario as mean % $\pm$ standard deviation. In the temporal case, context signals are provided to the contextual integration network as GRU hidden state representations, conditioned on previous inputs in the sequence. Results are expressed for each scenario as mean % $\pm$ standard deviation with 3 random seeds.
@@ 3 Discussion
The ability to appropriately use and switch contexts to make sense of the vast stream of sensory information is an important feature of robust cognitive systems. It is thought that in the neocortex, contextual representations in higher order regions modulate lower sensory regions through top-down interactions arriving on the apical dendrites of pyramidal neurons. While it is known that apical dendrites modulate neuronal firing, their specific computational role in cognitive processes such as contextual integration remains poorly understood.
Here, we tested whether the apical compartment represents an architectural bias that is functionally useful for contextual integration in neural networks. To this end, we implemented (1) a neural network architecture with distinct apical compartments and a neuron-wise integration rule based on the observation that apical dendrites are gain modulators of somatic activity (Waters et al., 2003; Larkum et al., 2007), and developed (2) an ambiguous image classification task that requires contextual integration to be solved. We trained (1) the contextual modulator (top-down synapses onto apical dendrites) by gradient descent to simultaneously learn all input-context conditions defined in our task. We found that this bio-inspired dendritic architecture outperformed a single-compartment model which processes both sensory and contextual information in the compartment. In this task setting, we find there is an inherent benefit to have contextual information propagated via a distinct apical compartment.
By applying LRP, we find that the learned top-down modulation of pyramidal neurons is consistent with the principles of biased competition (Spratling, 2008). Specifically, by analyzing the apical amplitude on class-specific neurons for ambiguous stimuli, we found the highest gain modulation in class-specific neurons that matched the target class. Moreover, we found that the level of competition introduced by the context adapted to the level of ambiguity in the task: The top-down network learned to a) refine ambiguous representations by applying strong but sparse gain modulation to basal signals when contextual information is both available and relevant, and b) provide weak modulation in the case of contradictory or irrelevant contexts or unambiguous stimuli.
3.1 Comparison with existing computational frameworks
From a computational perspective, apical dendrites have been previously shown to support some types of contextual integration in biophysically realistic neurons and ANNs (Naumann et al., 2022; Wybo et al., 2022; George et al., 2020; Baronig and Legenstein, 2024). The tasks used in these previous works are multi-task, single scenario problem settings, meaning that context (e.g. one-hot encoded task id) is always assumed to be relevant. Alternative approaches to flexible contextual integration have involved explicitly learning to prevent associations between irrelevant context and sensory inputs ((Baronig and Legenstein, 2024)), which represents a successful example of how to handle spurious contexts in a simple task. As an extension of this approach, our model must infer how to handle the interaction between sensory input and context across distinct tasks and scenarios, where context is relevant, irrelevant, or contradictory.
Crucially, we only found a significant advantage of the apical prior over other integration approaches in the scenario where inputs are unambiguous and context is contradictory (and should be ignored), with an accuracy increase of 1.9% (MNIST) for models with distinct apical compartments. This highlights a specific advantage for apical dendrites in the role of contextual integration, where contextual relevance is crucial to performance. Furthermore, while the task can be solved using either the Hadamard or additive rule (Appendix 4), we found that, under uncertainty, the Hadamard integration rule enabled the model to ignore contradictory contextual signals more effectively compared to additive integration.
Another key difference with most previous work is that as basal and apical signals are combined post-activation, our rule can only amplify already active neurons, unlike Wybo et al. (2022) and (Naumann et al., 2022), but inline with experimental evidence (Larkum et al., 1999). From an efficiency perspective, this maintains the sparsity of the network.
One notable prior work which applies context-dependent modulation to feedforward networks is context-dependent gating (XdG) (Masse et al., 2018), which could be comparable with our basal-only ablation as it is a feedforward architecture. The main difference with our basal-only ablation is that XdG a) is not informed by intermediate sensory representations, and as most prior works, assumes context always agrees with sensory input. Therefore, its is unlikely that XdG, implemented as-is, can deal with cases where context should be overridden by sensory input, as in the unambig.Contra case.
The advantage of apical dendrites are clear in our multi-scenario contextual integration task. In order to compare our results with previous work, we examined the effect of our integration mechanism in a standard multi-task setting (given task-ids). Under these conditions we found that that all variants perform similarly well. This can be explained by the setting being simpler, whereas the distinct apical compartment shines under more varied and challenging types of contextual integration that require inference of ambiguity.
While not explicitly modular, our framework could be seen as in accordance with the Global Workspace Theory along with the "information bottleneck" (to discard irrelevant information, which can include contextual information) Goyal et al. (2022). Under both frameworks, representations can be learned in a distributed and scalable manner, and mapped to local populations as context signals (in our case via apical compartments), and through biased competition, integrated with bottom-up representations to solve perceptual tasks. As highlighted in their work, this frees up the capacity of a learning system for more tasks, as it can more efficiently reuse what it has previously learned and identify task-relevant information, compared to single-compartment, single-integration models.
3.2 Predictions: What is necessary for robust contextual integration?
Until now, we have shown that in our contextual integration task, the apical prior provides a specific advantage compared to single-compartment integration when sensory or contextual signals are unreliable. Here, we aim to extend these findings to make broader predictions for what specific components/computations might be necessary for robust contextual integration, and how such a mechanism could plausibly be implemented in the cortex.
Our results show that, given both sensory input and contextual input must be considered simultaneously, it may be sub-optimal to sum all activity together in the basal compartment, when sensory information and context are contradictory. At a higher level, this highlights a key difference between single and two-zone integration of information. As such, an interesting experiment would be to compare the single vs two-zone integration of bottomup representations and topdown contextual representations in simulations of neurons which more faithfully respect the biophysical constraints of pyramidal neurons in the neocortex, to observe whether our results hold in a more biologically plausible setting, or if it is strictly a feature of ANNs.
Furthermore, based on our results, we can infer that the apical compartment encodes some information about the ambiguity of the sensory input, and the relevance of context for a given sensory input. An interesting experiment could involve silencing or deactivating apical dendrites in more realistic neurons (biophysical simulation or biological), and observing the effect on accurate perception in a similar contextual integration task.
Mutual information between apical and basal activation is positive for the ambig.Match scenario, where we see a reduction in entropy of the prediction after observing the contextual signal. On the other hand, mutual information is near 0 for the unambig.Contra scenario, which aligns with our main findings (see Appendix, Table 3). This is not surprising given that we performed supervised training. We explicitly trained under the assumption that local representations, given mis-matching context, should be invariant to sensory ambiguity. This could be interpreted as being surprised by a stimulus, but not having informative or relevant context to reduce the ambiguity of the stimulus. This is important for robust contextual integration if one subscribes to the hypothesis that, most available contextual signals in rich, real-world sensory experience are distractors for any one particular task, and should be discarded.
Following this, we predict that a model cannot learn the relevance of context with out encountering scenarios with irrelevant context during training. As an ablation study, we could remove the unambig.Irrel scenario from our loss function. We expect to observe a significant drop in accuracy on that scenario, reflecting the inability to assess contextual relevance.
Extending this idea to unsupervised learning, it is interesting to think about how the model could identify contextual relevance without a supervised signal and be invariant to sensory ambiguity under unhelpful context. In an unsupervised setting this could arise from an information-maximization learning rule like InfoNCE, which uses a contrastive loss with positive and negative pairs defined via augmentation of the input (positive) against other inputs (negative). Unfortunately, the creation of positive and negative pairs relies heavily on domain-specific knowledge for useful data augmentations (eg. a rotation applied on an image). If it is possible to create contrastive pairs in a domain-agnostic manner, such an unsupervised learning framework could approximate the effects of supervised training shown in our results.
Given the fact that most associations of context and input are either irrelevant or contradictory, and only a small subset of associations represent matching context, it is not surprising that the gain modulation is sparse. We expect also that our high-dimensional apical weight matrices could be well approximated with low-rank matrices without losing significant performance, in line with work in modulating deep learning models for downstream tasks via low-rank adapters Hu et al. (2021). While this method uses resources efficiently, it lacks the mechanism to separate irrelevant or contradictory from useful contexts that the apical prior excels at, which indicates a potential synergy between the two. We hypothesize the brain uses both mechanisms to learn to handle the large pool of available sources of context efficiently. A future direction could combine LORAs with the apical prior for contextual integration tasks.
We find that our model could learn a mapping of top-down signals onto the apical compartment, which importantly, was learned separately from the contextual representations. This is compatible with high-order cortical regions such as the hippocampus, where the learning and updating of contextual representations is disentangled from how they modulate sensory representations (Maren et al., 2013). Because we assume the stability of the sensory and contextual representation, we cannot make any conclusions about how these two representations could be simultaneously co-learned. However, we speculate there should be some degree of stability in the sensory representations to learn to perform contextual integration.
3.3 Limitations
While our model unifies the function of contextual integration with key biophysical properties of dendrites, some key model assumptions and simplifications are worth discussing with respect to their empirical support.
@@ 3.3 Limitations
Firstly, our model is designed to implement some key features of cortical pyramidal neurons without focusing on a specific subpopulation that make up cortical layers, including layer 2/3 and layer 5 pyramidal neurons. Here, we assume that contextual signals are provided by neurons in a higher-order region, which would likely correspond to layer 5 output, and integrated in apical dendrites of putative layer 2/3 pyramidal neurons (Bastos et al., 2012; Schuman et al., 2021). It is noteworthy though that layer 5 neurons can also send top-down projections to the basal compartment of layer 2/3 neurons, although to a lesser extent (Harris et al., 2019). Unlike other computational frameworks such as (George et al., 2020) which seek to incorporate the rich architectural diversity of the neocortex, we sought to simplify and abstract many features (using neural networks) in order to be able to train models to solve more challenging contextual integration tasks which could give useful insights that could not be seen with simpler tasks.
@@
Our model heavily abstracts temporal dynamics that are normally observed in biological circuits. For example, in the presence of ambiguous stimuli, human subjects can oscillate between two interpretations, a phenomenon termed ’perceptual bistability’ (Hardstone et al., 2021). During perceptual bistability, neural activity alternates between discrete states (Abeles et al., 1995; Akrami et al., 2009), following principles of continuous attractor dynamics (Daelli and Treves, 2010).
3.4 Outlook
In the future, our model could support more realistic temporal dynamics, which would enable the testing of a rich array of hypotheses regarding contextual integration.
In the future, we propose that our model could be extended to hierarchies of sensory modalities interacting with each other. Deep learning models that implement multi-modal interactions are an active area of research (Radford et al., 2021; Alayrac et al., 2022; Mustafa et al., 2022). While they are very good at multimodal generation with conditioning on one modality, they do not currently have the ability to resolve conflicting information and can be easily confused by contradictions in mixed-modality data. While these models employ transformer architectures that integrate context differently than MLPs, our loss function could inspire new approaches for robust multi-modal perception in deep neural networks.
From a neurophysiology perspective, our framework could be leveraged to test specific hypotheses of contextual integration in health and disease, and in particular in cases where the physiological properties of pyramidal neurons are known to be altered (Kim et al., 2014).